# thanks: M.C. and D.P. contributed equally to this work

### v1

@@ II.1 Simpson-Visser space-time
The Simpson-Visser (SV) metric is a one-parameter extension of the Schwarzschild space-time and easily one of the best known black-bounce metrics. In the words of Simpson and Visser, in some sense this metric “represents the minimal violence to the standard Schwarzschild solution” needed to enforce regularity Simpson and Visser (2019). The line element analytically interpolates between black holes and traversable wormhole according to the value of the regularizing parameter. In the notation of Eq. (1), the SV space-time is characterized by the following functions:  [^3] For all the space-times we will consider, the parameter $M$ appearing in the function $f(r)$ always corresponds to the mass of the space-time (either the Komar, ADM, Misner-Sharp-Hernandez, or Brown-York mass).
@@ II.3 D’Ambrosio-Rovelli space-time
The D’Ambrosio-Rovelli (DR) space-time was originally developed with motivations other than singularity avoidance, and is in fact also somewhat motivated by LQG considerations. This space-time represents a natural extension of the Schwarzschild space-time which crosses the $r=0$ singularity smoothly into the interior of a white hole, and one can see it as the $\hbar\to 0$ limit of an effective QG metric Bianchi et al. (2018); D’Ambrosio and Rovelli (2018). It has been argued that this black hole-to-white hole tunneling mechanism can shed light on possible solutions to the information paradox. Of interest to us is the fact that the DR metric is regular, as a result of the curvature of the effective metric being bound a the Planck scale. The ansatz for the effective metric written by D’Ambrosio and Rovelli is similar to that of the SV metric, but differs in terms of the $\mathfrak{g}$ function – specifically, the two relevant functions are given by the following:
@@ III.1 Greybody factors
We now discuss the computation of the greybody factors (GBFs), functions of energy and angular momentum which characterize the shape of the emitted Hawking radiation (and in particular its deviation from a blackbody) and therefore play a key role in determining the resulting evaporation constraints Sakalli and Kanzi (2022); Konoplya and Zhidenko (2024b, c). It is worth noting that, in the notation of Eq. (3), both the SV and PK metrics share the fact that $f(r)=g(r)$, which makes the calculations significantly easier. This is not the case, however, for the DR space-time. Therefore, in what follows, we consider the more general case where $f(r)\neq g(r)$, which also differs from the simpler tr-symmetric case considered in our companion paper Calzà et al. (tion).
We adopt the Newman-Penrose (NP) formalism Newman and Penrose (1962), denoting by $\Upsilon_{s}$ a general perturbation of spin $s$ (defined by the appropriate NP scalars) and dropping the $l$ and $m$ indices to lighten the notation. Then, the Teukolsky equation for the evolution of massless perturbations of given spin upon a background metric characterized by functions $f(r)$, $g(r)$, $h(r)$ as in Eq. (3) reduces to the following master equation Teukolsky (1973):
@@
as proven in more detail in Appendix A (the case of the DR metric is not trivial).
To compute the GBFs, we make use of the shooting method, widely used earlier in similar contexts (see e.g. Refs. Rosa (2017, 2013); Calzà et al. (2021); Calzá (2023); Calzà and Rosa (2022); Calzà et al. (2023); Calzà and Rosa (2023a, b)), including in our companion paper Calzà et al. (tion). We begin by defining the rescaled coordinate $x$:
@@
$\left.\left.\left.-s\dot{g}\left(2i\omega\sqrt{\frac{f}{g}+\dot{f}}\right)\right)\right)+sh\left((3-2s)\dot{f}+4i\omega\sqrt{\frac{f}{g}}\right)\right)\,.$
while the dot denotes a derivative with respect to the rescaled coordinate $x$. For completeness, we note that $f$ and $g$ as as function of $x$ for the metrics considered here are given by the following:
$\begin{split}&h_{\text{SV}}(x)=h_{\text{PK}}(x)=h_{\text{DR}}(x)=(x+1)^{2}+\ell^{2}\\ &g_{\text{SV}}(x)=f_{SV}(x)=1-\frac{\sqrt{1+\ell^{2}}}{\sqrt{\ell^{2}+(x+1)^{2}}}\\ &g_{\text{PK}}(x)=f_{\text{PK}}(x)=\frac{x}{\sqrt{\ell^{2}+(x+1)^{2}}}\\ &g_{\text{DR}}(x)=1-\frac{\sqrt{1+\ell^{2}}}{\sqrt{\ell^{2}+(x+1)^{2}}}\\ &f_{\text{DR}}(x)=g_{\text{DR}}(x)\left(1+\frac{\ell}{\sqrt{\ell^{2}+(x+1)^{2}}}\right)^{-1}\,.\end{split}$ (26)
@@
Here, $\tau$ is a function of the field’s spin and regularizing parameter $\ell$, and varies with the metric being considered. For further details, we refer the reader to the Appendix of our companion paper Calzà et al. (tion), where the issue is discussed in more depth. We then determine the coefficients $a_{n}$ iteratively, by repeatedly substituting Eq. (27) into Eq. (22).
@@
As can be seen in Eqs. (31,32), deviations from $\alpha=1$ are associated to the non-tr-symmetric nature of the metrics. As a general consideration, we note that the computation of GBFs for the metrics under consideration is more involved compared to those considered in our companion paper Calzà et al. (tion).
@@ III.2 Evaporation spectra
As in our companion paper Calzà et al. (tion), we only consider the primary photon spectrum resulting from Hawking evaporation, while also checking that within the mass region of interest the secondary spectrum will provide a negligible contribution. For a particle species $i$ of spin $s$ (characterized by $n_{i}$ degrees of freedom), the rate of emission (particles per unit time per unit energy) through Hawking radiation is given by the following Hawking (1975a, b); Page (1976a, b, 1977):
@@ III.3 Evaporation constraints
In the mass range $10^{10}\,{\text{kg}}\lesssim M_{\text{pbh}}\lesssim 10^{15}\,{\text{kg}}$, the dominant constraints on the PBH abundance come from measurements of the extragalactic photon background Auffinger (2023), and more precisely of the diffuse extragalactic $\gamma$-ray background (EGRB) in the energy range $100\,{\text{keV}}\lesssim E_{\gamma}\lesssim 5\,{\text{GeV}}$, which can be directly compared against theoretical expectations for their Hawking evaporation spectra. Of particular interest to us is the fact that the lower edge of the asteroid mass window, where PBHs could make up the entire DM, is set precisely by evaporation constraints. In what follows, we set evaporation constraints on the fraction of DM in the form of PBHs $f_{\text{pbh}}(M)\equiv\Omega_{\text{pbh}}/\Omega_{\text{dm}}$, assuming that PBHs are described by the three metrics discussed so far. We work under the same set of approximations adopted in our companion paper Calzà et al. (tion). Namely, we assume that PBHs are isotropically distributed on sufficiently large scales and cluster in the galactic halo in the same way as other forms as DM, we only compute the primary photon spectrum, and finally we assume a monochromatic mass distribution (see e.g. Refs. Kühnel et al. (2016); Kühnel and Freese (2017); Carr et al. (2017); Raidal et al. (2017); Bellomo et al. (2018); Lehmann et al. (2018); Carr and Kuhnel (2019); Gow et al. (2020); De Luca et al. (2020d); Gow et al. (2022); Ashoorioon et al. (2021); Bagui and Clesse (2022); Mukhopadhyay et al. (2022); Papanikolaou (2022); Cai et al. (2024) for studies on the effects of an extended mass distribution). We refer the reader to Sec. IIIC of our companion paper Calzà et al. (tion) for a more detailed discussion of why these, which are clearly all approximations, are appropriate for the scope of our work (while allowing for a more direct comparison to earlier works, including our companion paper).
@@
In Eqs. (35,36), $d^{2}{N}_{\gamma}/dtdE_{\gamma}$ is computed via Eq. (34), whereas $z_{\star}$ is the redshift of recombination, $H(z)$ is the expansion rate, and $n_{\text{pbh}}(t_{0})$ is the present-day PBH number density, itself related to the parameter of interest $f_{\text{pbh}}$ via the following:
@@ IV Results
Figure 5: Upper limits on $f_{\text{pbh}}$, the fraction of dark matter in the form of primordial regular Simpson-Visser black holes, as a function of the black hole mass $M_{\text{pbh}}$. The limits are derived for different values of the regularizing parameter $\ell$ (normalized by the horizon radius $r_{H}$), with the shaded regions excluded: $\ell/r_{H}=0.3$ (red dotted curve), $0.6$ (green dashed curve), and $0.9$ (magenta dash-dotted curve). Note that the blue solid curve corresponds to the case $\ell/r_{H}=0$, which recovers the Schwarzschild black hole, whereas the value of $M_{\text{pbh}}$ corresponding to the upper right edge of the $f_{\text{pbh}}$ constraints marks the lower edge of the asteroid mass window.
@@
The resulting upper limits on $f_{\text{pbh}}$ as a function of PRBH mass $M_{\text{pbh}}$, for different values of $\ell$, are shown in Figs. 5,  6, and 7 for the Simpson-Visser, Peltola-Kunstatter, and D’Ambrosio-Rovelli space-times respectively. In each figure, shown as blue solid curves are the corresponding constraints in the Schwarzschild PBH case ($\ell\to 0$), which we have verified to recover the results of Ref. Carr et al. (2010). We stress that for a given value of $\ell$, the value of $M_{\text{pbh}}$ where the overclosure limit $f_{\text{pbh}}<1$ is saturated sets the lower edge of the modified asteroid mass window (potentially enlarged or contracted).
@@
For all three cases, we see that increasing the regularizing parameter $\ell$ results (at a given value of $M_{\text{pbh}}$) in weaker limits on $f_{\text{pbh}}$. This confirms the expectation raised at the end of Sec. III.2 upon inspection of the resulting evaporation spectra, all of which decrease in intensity relative to the Schwarzschild case (except for the slight increase in the Peltola-Kunstatter BH case for energies below the peak, which we recall reflects the different behaviour of the GBFs). We see that, for $\ell/r_{H}=0.9$, the upper limits on $f_{\text{pbh}}$ weaken by up to an order of magnitude at a given $M_{\text{pbh}}$ relative to the Schwarzschild limit for the Simpson-Visser and D’Ambrosio-Rovelli BHs, whereas for the Peltola-Kunstatter BH the extent to which the $f_{\text{pbh}}$ limits weaken is more limited (again unsurprisingly, given the enhanced intensity of the evaporation spectrum at low energies).
@@
The aforementioned shifts result in the asteroid mass window being enlarged for all three metrics considered, because the lower edge of the window (lying roughly at $M_{\text{pbh}}\simeq 10^{14}\,{\text{kg}}$ in the Schwarzschild case) moves towards lower masses. In general, we observe that the asteroid mass window further opens up by about half a decade in mass (an increase which is less dramatic than what we observed for the phenomenological metrics in our companion paper Calzà et al. (tion)). As a result, there is a wider available range of parameter space where PBHs of the type we are considering could make up all the DM. We note that our constraints assume that all PRBHs in the Universe carry the same value of “hair parameter” $\ell/r_{H}$ (in the language of Ref. Vagnozzi et al. (2023), we are treating it as an “universal hair”). Whether or not this is a reasonable assumption requires a deeper investigation of the theoretical underpinning of the adopted metrics (in particular the LQG-inspired one) which, in the spirit of the present work being a pilot study, we defer to follow-up work.
@@
Finally, in our companion paper Calzà et al. (tion) we extensively commented on a few caveats concerning the extension of the asteroid mass window and, more generally, on other existing constraints on $f_{\text{pbh}}$, which bear repeating here, albeit in a more condensed form (we refer the reader to Sec. IV of our companion paper for a significantly more detailed discussion). While evaporation constraints set the lower edge of the asteroid mass window, the upper edge thereof is instead set by lensing constraints. Our claim that the asteroid mass window is enlarged because the lower edge moves towards even lower values is therefore contingent upon the upper edge remaining the standard Schwarzschild one, even within the adopted metrics. We do, in fact, expect this to be the case since, at fixed $M_{\text{pbh}}$, lensing constraints depend only on the mass of the lensing object. We can therefore assert that the asteroid mass window is indeed enlarged when considering the three PBH metrics introduced here. Finally, a variety of other constraints on $f_{\text{pbh}}$ exist, including dynamical, accretion, and CMB constraints (see e.g. Ref. Carr and Kuhnel (2022) for a recent summary): however, with the exception of a few debated constraints Capela et al. (2013a, b); Pani and Loeb (2014); Graham et al. (2015), we expect these to be relevant within significantly different mass ranges (unless accretion dynamics are significantly different around the RBHs under consideration), although we reserve a detailed study to follow-up work.
@@ V Conclusions
It is a commonly held belief that the singularities which plague General Relativity, and represent one of the most important open problems in theoretical physics, will eventually be solved once the long sought after theory of quantum gravity is unveiled. While a consensus theory of quantum gravity remains elusive, progress on the singularity problem can still be made by considering ansätze for singularity-free space-times, either introduced phenomenologically or somewhat motivated from candidate quantum gravity frameworks (such as LQG). These regular black holes, if produced early on in the Universe from the collapse of large density perturbations (thus being primordial regular BHs), could also have a role to play in the dark matter problem. Our work is a pilot study which goes precisely in this direction, examining what are the consequences of PBHs being described by non-singular metrics. In fact, it bears reminding that the usual constraints on the fraction of DM in the form of PBHs, $f_{\text{pbh}}$, are derived under the assumption of PBHs being described by the Schwarzschild metric, which is well-known to be plagued by the $r=0$ singularity. In the present work, we have explored three so-called non-tr-symmetric metrics as candidates for describing PBHs: the Simpson-Visser black-bounce, Peltola-Kunstatter, and D’Ambrosio-Rovelli black-to-white-hole-bounce space-times, with the latter two enjoying strong theoretical motivation from LQG (we note that the mathematically simpler tr-symmetric case is covered in our companion paper Calzà et al. (tion), and includes well-known phenomenological regular BHs such as the Bardeen and Hayward BHs).
@@
After discussing the impact of the regularizing parameter $\ell$ (with the Schwarzschild BH corresponding to the $\ell\to 0$ limit) on the resulting evaporation spectra, we find that as $\ell$ increases, at a fixed PBH mass $M_{\text{pbh}}$ the corresponding upper limit on $f_{\text{pbh}}$ weakens for all three metrics considered. This results in the lower edge of the asteroid mass window shifting down by up to approximately half a decade in $M_{\text{pbh}}$ parameter space (down from $M_{\text{pbh}}\simeq 10^{14}\,{\text{kg}}$ to $M_{\text{pbh}}\simeq 3\times 10^{13}\,{\text{kg}}$). As a result, there is a larger range of available parameter space where the PBHs in question could make up the entire DM in the Universe, which could be targeted by proposed probes of the asteroid mass window Ray et al. (2021); Ghosh and Mishra (2024); Branco et al. (2023); Bertrand et al. (2023); Tran et al. (2023); Dent et al. (2024); Tamta et al. (2024).
@@
Our work (alongside our companion paper Calzà et al. (tion)) demonstrates, as a proof-of-principle, that the intersection of the DM and singularity problems is a fertile terrain worthy of further studies. The most important avenue for immediate follow-up work would be to systematically revisit, within the metrics under consideration, all other non-evaporation constraints which have been extensively discussed for Schwarzschild PBHs (including lensing, accretion, and dynamical constraints). Moreover, since the Peltola-Kunstatter and D’Ambrosio-Rovelli space-times are rooted within an underlying quantum gravity theoretical framework, a first-principles study of their formation mechanism (which is otherwise not possible for metrics introduced at a phenomenological level, such as the Bardeen and Hayward BHs) and whether this leads to additional interesting complementary signatures is definitely worth pursuing. For instance, quantum transitions near the would-be singularity of the D’Ambrosio-Rovelli space-time should lead to the existence of long-lived (primordial) white holes, which in turn could potentially lead to a wide range of exotic signatures one could hope to search for. In continuing our exciting program at the interface of the DM and singularity problems, it is our intention to return to these and related issues in future follow-up work.

### v2

@@ II.1 Simpson-Visser space-time
The Simpson-Visser (SV) metric is a one-parameter extension of the Schwarzschild space-time and easily one of the best known black-bounce metrics. In the words of Simpson and Visser, in some sense this metric “represents the minimal violence to the standard Schwarzschild solution” needed to enforce regularity Simpson and Visser (2019). The line element analytically interpolates between black holes and traversable wormholes according to the value of the regularizing parameter. In the notation of Eq. (1), the SV space-time is characterized by the following functions:  [^3] For all the space-times we will consider, the parameter $M$ appearing in the function $f(r)$ always corresponds to the mass of the space-time (either the Komar, ADM, Misner-Sharp-Hernandez, or Brown-York mass).
@@ II.3 D’Ambrosio-Rovelli space-time
The D’Ambrosio-Rovelli (DR) space-time was originally developed with motivations other than singularity avoidance, and is in fact also motivated by LQG considerations. This space-time represents a natural extension of the Schwarzschild space-time which crosses the $r=0$ singularity smoothly into the interior of a white hole, and one can see it as the $\hbar\to 0$ limit of an effective QG metric Bianchi et al. (2018); D’Ambrosio and Rovelli (2018). It has been argued that this black hole-to-white hole tunneling mechanism can shed light on possible solutions to the information paradox. Of interest to us is the fact that the DR metric is regular, as a result of the curvature of the effective metric being bound at the Planck scale. The ansatz for the effective metric written by D’Ambrosio and Rovelli is similar to that of the SV metric, but differs in the form of the $\mathfrak{g}$ function – specifically, the two relevant functions are given by the following:
@@ III.1 Greybody factors
We now discuss the computation of the greybody factors (GBFs), functions of energy and angular momentum which characterize the shape of the emitted Hawking radiation (and in particular its deviation from a blackbody) and therefore play a key role in determining the resulting evaporation constraints Sakalli and Kanzi (2022); Konoplya and Zhidenko (2024b, c). It is worth noting that, in the notation of Eq. (3), both the SV and PK metrics share the fact that $f(r)=g(r)$, which makes the calculations somewhat easier. This is not the case, however, for the DR space-time. Therefore, in what follows, we consider the more general case where $f(r)\neq g(r)$, which differs from the much simpler tr-symmetric case considered in our companion paper Calzà et al. (2024a).
We adopt the Newman-Penrose (NP) formalism, denoting by $\Upsilon_{s}$ a general perturbation of spin $s$ (defined by the appropriate NP scalars) and dropping the $l$ and $m$ indices to lighten the notation. Then, the Teukolsky equation for the evolution of massless perturbations of given spin upon a background metric characterized by functions $f(r)$, $g(r)$, $h(r)$ as in Eq. (3) reduces to the following master equation Teukolsky (1973):
@@
as proven in more detail in Appendix A (the case of the DR metric is actually far from trivial).
To compute the GBFs, we make use of the shooting method, widely used earlier in similar contexts (see e.g. Refs. Rosa (2017, 2013); Calzà et al. (2021); Calzá (2023); Calzà and Rosa (2022); Calzà et al. (2023); Calzà and Rosa (2023a, b)), including in our companion paper Calzà et al. (2024a). We begin by defining the rescaled coordinate $x$:
@@
$\left.\left.\left.-s\dot{g}\left(2i\omega\sqrt{\frac{f}{g}+\dot{f}}\right)\right)\right)+sh\left((3-2s)\dot{f}+4i\omega\sqrt{\frac{f}{g}}\right)\right)\,,$
where the dot denotes a derivative with respect to the rescaled coordinate $x$. For completeness, we note that $f$ and $g$ as as function of $x$ for the metrics considered here are given by the following:
$\begin{split}&h_{\text{SV}}(x)=h_{\text{PK}}(x)=h_{\text{DR}}(x)=(x+1)^{2}+\ell^{2}\,,\\ &g_{\text{SV}}(x)=f_{SV}(x)=1-\frac{\sqrt{1+\ell^{2}}}{\sqrt{\ell^{2}+(x+1)^{2}}}\,,\\ &g_{\text{PK}}(x)=f_{\text{PK}}(x)=\frac{x}{\sqrt{\ell^{2}+(x+1)^{2}}}\,,\\ &g_{\text{DR}}(x)=1-\frac{\sqrt{1+\ell^{2}}}{\sqrt{\ell^{2}+(x+1)^{2}}}\,,\\ &f_{\text{DR}}(x)=g_{\text{DR}}(x)\left(1+\frac{\ell}{\sqrt{\ell^{2}+(x+1)^{2}}}\right)^{-1}\,.\end{split}$ (26)
@@
Here, $\tau$ is a function of the field’s spin and regularizing parameter $\ell$, and varies with the metric being considered. For further details, we refer the reader to the Appendix of our companion paper Calzà et al. (2024a), where the issue is discussed in more depth. We then determine the coefficients $a_{n}$ iteratively, by repeatedly substituting Eq. (27) into Eq. (22).
@@
As can be seen in Eqs. (31,32), deviations from $\alpha=1$ are associated to the non-tr-symmetric nature of the metrics. As a general consideration, we note that the computation of GBFs for the metrics under consideration is significantly more involved compared to those considered in our companion paper Calzà et al. (2024a).
@@ III.2 Evaporation spectra
As in our companion paper Calzà et al. (2024a), we only consider the primary photon spectrum resulting from Hawking evaporation, while also checking that within the mass region of interest the secondary spectrum will provide a negligible contribution. For a particle species $i$ of spin $s$ (characterized by $n_{i}$ degrees of freedom), the rate of emission (particles per unit time per unit energy) through Hawking radiation is given by the following Hawking (1975a, b); Page (1976a, b, 1977):
@@ III.3 Evaporation constraints
In the mass range $10^{10}\,{\text{kg}}\lesssim M_{\text{pbh}}\lesssim 10^{15}\,{\text{kg}}$, the dominant constraints on the PBH abundance come from measurements of the extragalactic photon background Auffinger (2023), and more precisely of the diffuse extragalactic $\gamma$-ray background (EGRB) in the energy range $100\,{\text{keV}}\lesssim E_{\gamma}\lesssim 5\,{\text{GeV}}$, which can be directly compared against theoretical expectations for the PBH Hawking evaporation spectra. Of particular interest to us is the fact that the lower edge of the asteroid mass window, where PBHs could make up the entire DM, is precisely set by evaporation constraints. In what follows, we set evaporation constraints on the fraction of DM in the form of PBHs $f_{\text{pbh}}(M)\equiv\Omega_{\text{pbh}}/\Omega_{\text{dm}}$, assuming that PRBHs are described by the three metrics discussed so far. We work under the same set of approximations adopted in our companion paper Calzà et al. (2024a). Namely, we assume that PRBHs are isotropically distributed on sufficiently large scales and cluster in the galactic halo in the same way as other forms as DM, we only compute the primary photon spectrum, and finally we assume a monochromatic mass distribution (see e.g. Refs. Kühnel et al. (2016); Kühnel and Freese (2017); Carr et al. (2017); Raidal et al. (2017); Bellomo et al. (2018); Lehmann et al. (2018); Carr and Kuhnel (2019); Gow et al. (2020); De Luca et al. (2020d); Gow et al. (2022); Ashoorioon et al. (2021); Bagui and Clesse (2022); Mukhopadhyay et al. (2022); Papanikolaou (2022); Cai et al. (2024b) for studies on the effects of an extended mass distribution). We refer the reader to Sec. IIIC of our companion paper Calzà et al. (2024a) for a more detailed discussion of why these, which are clearly all approximations, are appropriate for the scope of our work (while allowing for a more direct comparison to earlier works, including our companion paper).
@@
In Eqs. (35,36), $d^{2}{N}_{\gamma}/dtdE_{\gamma}$ is computed via Eq. (34), whereas $z_{\star}$ is the redshift of recombination, $H(z)$ is the expansion rate, and $n_{\text{pbh}}(t_{0})$ is the present-day PRBH number density, itself related to the parameter of interest $f_{\text{pbh}}$ via the following:
@@ IV Results
Figure 5: Upper limits on $f_{\text{pbh}}$, the fraction of dark matter in the form of primordial regular Simpson-Visser black holes, as a function of the black hole mass $M_{\text{pbh}}$. The limits are derived for different values of the regularizing parameter $\ell$ (normalized to the horizon radius $r_{H}$), with the shaded regions excluded: $\ell/r_{H}=0.3$ (red dotted curve), $0.6$ (green dashed curve), and $0.9$ (magenta dash-dotted curve). Note that the blue solid curve corresponds to the case $\ell/r_{H}=0$, which recovers the Schwarzschild black hole, whereas the value of $M_{\text{pbh}}$ corresponding to the upper right edge of the $f_{\text{pbh}}$ constraints marks the lower edge of the asteroid mass window.
@@
The resulting upper limits on $f_{\text{pbh}}$ as a function of PRBH mass $M_{\text{pbh}}$, for different values of $\ell$, are shown in Figs. 5,  6, and 7 for the SV, PK, and DR space-times respectively. In each figure, shown as blue solid curves are the corresponding constraints in the Schwarzschild PBH case ($\ell\to 0$), which we have verified to recover the results of Ref. Carr et al. (2010). We stress that for a given value of $\ell$, the value of $M_{\text{pbh}}$ where the overclosure limit $f_{\text{pbh}}<1$ is saturated sets the lower edge of the modified asteroid mass window (potentially enlarged or contracted).
@@
For all three cases, we see that increasing the regularizing parameter $\ell$ results (at a given value of $M_{\text{pbh}}$) in weaker limits on $f_{\text{pbh}}$. This confirms the expectation raised at the end of Sec. III.2 upon inspection of the resulting evaporation spectra, all of which decrease in intensity relative to the Schwarzschild case (except for the slight increase in the PK PRBH case for energies below the peak, which we recall reflects the different behaviour of the GBFs). We see that, for $\ell/r_{H}=0.9$, the upper limits on $f_{\text{pbh}}$ weaken by up to an order of magnitude at a given $M_{\text{pbh}}$ relative to the Schwarzschild limit for the SV and DR PRBHs, whereas for PK PRBHs the extent to which the $f_{\text{pbh}}$ limits weaken is more limited – again unsurprisingly, given that the enhanced intensity of the evaporation spectrum at low energies counteracts the decrease at higher energies in the integral of Eq. (35).
@@
The aforementioned shifts result in the asteroid mass window being enlarged for all three metrics considered, because the lower edge of the window (lying roughly at $M_{\text{pbh}}\simeq 10^{14}\,{\text{kg}}$ in the Schwarzschild case) moves towards lower masses. In general, we observe that the asteroid mass window further opens up by about half a decade in mass or more (an increase which is less dramatic than what we observed for the phenomenological metrics in our companion paper Calzà et al. (2024a)). As a result, there is a wider available range of parameter space where PRBHs of the type we are considering could make up all the DM. We note that our constraints assume that all PRBHs in the Universe carry the same value of “hair parameter” $\ell/r_{H}$ (in the language of Ref. Vagnozzi et al. (2023), we are treating it as an “universal hair”). Whether or not this is a reasonable assumption requires a deeper investigation of the theoretical underpinning of the adopted metrics (in particular the LQG-inspired one) which, in the spirit of the present work being a pilot study, we defer to follow-up work.
@@
Finally, in our companion paper Calzà et al. (2024a) we extensively commented on a few caveats concerning the extension of the asteroid mass window and, more generally, on other existing constraints on $f_{\text{pbh}}$, which bear repeating here, albeit in a more condensed form (we refer the reader to Sec. IV of our companion paper for a significantly more detailed discussion). While evaporation constraints set the lower edge of the asteroid mass window, the upper edge thereof is instead set by lensing constraints. Our claim that the asteroid mass window is enlarged because the lower edge moves towards even lower values is therefore contingent upon the upper edge remaining the standard Schwarzschild one, even within the adopted metrics. We do, in fact, expect this to be the case since, at fixed $M_{\text{pbh}}$, lensing constraints depend only on the mass of the lensing object. We can therefore assert that the asteroid mass window is indeed enlarged when considering the three PBH metrics introduced here. Finally, a variety of other constraints on $f_{\text{pbh}}$ exist, including dynamical, accretion, and CMB constraints (see e.g. Ref. Carr and Kuhnel (2022) for a recent summary): however, with the exception of a few debated constraints Capela et al. (2013a, b); Pani and Loeb (2014); Graham et al. (2015), we expect these to be relevant within significantly different mass ranges (unless accretion dynamics are significantly different around the RBHs under consideration), although we reserve a detailed study to follow-up work.
@@ V Conclusions
It is a commonly held belief that the singularities which plague General Relativity, and represent one of the most important open problems in theoretical physics, will eventually be solved once the long sought after theory of quantum gravity is unveiled. While a consensus theory of quantum gravity remains elusive, progress on the singularity problem can still be made by considering ansätze for singularity-free space-times, either introduced phenomenologically or somewhat motivated from candidate quantum gravity frameworks (such as LQG). These regular black holes, if produced early on in the Universe from the collapse of large density perturbations (thus being primordial regular BHs), could also have a role to play in the dark matter problem. Our work is a pilot study which goes precisely in this direction, examining what are the consequences of PBHs being described by non-singular metrics. In fact, it bears reminding that the usual constraints on the fraction of DM in the form of PBHs, $f_{\text{pbh}}$, are derived under the assumption of PBHs being described by the Schwarzschild metric, which is well-known to be plagued by the $r=0$ singularity. In the present work, we have explored three so-called non-tr-symmetric metrics as candidates for describing PBHs: the Simpson-Visser black-bounce, Peltola-Kunstatter, and D’Ambrosio-Rovelli black-to-white-hole-bounce space-times, with the latter two enjoying strong theoretical motivation from LQG (we note that the mathematically simpler tr-symmetric case is covered in our companion paper Calzà et al. (2024a), and includes well-known phenomenological regular BHs such as the Bardeen and Hayward BHs).
@@
After discussing the impact of the regularizing parameter $\ell$ (with the Schwarzschild BH corresponding to the $\ell\to 0$ limit) on the resulting evaporation spectra, we find that as $\ell$ increases, at a fixed PRBH mass $M_{\text{pbh}}$ the corresponding upper limit on $f_{\text{pbh}}$ from observations of the EGRB weakens for all three metrics considered. This results in the lower edge of the asteroid mass window shifting down by up to approximately half a decade in $M_{\text{pbh}}$ parameter space (down from $M_{\text{pbh}}\simeq 10^{14}\,{\text{kg}}$ to $M_{\text{pbh}}\simeq 3\times 10^{13}\,{\text{kg}}$). As a result, there is a larger range of available parameter space where the PBHs in question could make up the entire DM in the Universe, which could be targeted by proposed probes of the asteroid mass window Ray et al. (2021); Ghosh and Mishra (2024); Branco et al. (2023); Bertrand et al. (2023); Tran et al. (2023); Dent et al. (2024); Tamta et al. (2024).
@@
Our work (alongside our companion paper Calzà et al. (2024a)) demonstrates, as a proof-of-principle, that the intersection of the DM and singularity problems is a fertile terrain worthy of further studies. The most important avenue for immediate follow-up work would be to systematically revisit, within the metrics under consideration, all other non-evaporation constraints which have been extensively discussed for Schwarzschild PBHs (including lensing, accretion, and dynamical constraints). Moreover, since the Peltola-Kunstatter and D’Ambrosio-Rovelli space-times are rooted within an underlying quantum gravity theoretical framework, a first-principles study of their formation mechanism (which is otherwise not possible for metrics introduced at a phenomenological level, such as the Bardeen and Hayward BHs) and whether this leads to additional interesting complementary signatures is definitely worth pursuing. For instance, quantum transitions near the would-be singularity of the D’Ambrosio-Rovelli space-time should lead to the existence of long-lived (primordial) white holes, which in turn could potentially lead to a wide range of exotic signatures one could hope to search for. In continuing our exciting program at the interface of the DM and singularity problems, it is our intention to return to these and related issues in future follow-up work.